# Dequantified Diffusion-Schrödinger Bridge for Density Ratio Estimation

### v3

@@ 4.1 Density Estimation
2-D Synthetic Datasets. We trained DRE-$\infty$ (baseline) and $\text{D}^{3}\text{RE}$ (ours) on eight 2-D synthetic datasets, including $\mathsf{swissroll}$, $\mathsf{circles}$, $\mathsf{rings}$, $\mathsf{moons}$, $\mathsf{8gaussians}$, $\mathsf{pinwheel}$, $\mathsf{2spirals}$, and $\mathsf{checkerboard}$, for 20,000 epochs using the joint score matching loss (details in Section C.2). The density estimation results are shown in Figure 2.
@@
In contrast, the $\text{D}^{3}\text{RE}$ framework, leveraging both DDBI and DSBI, achieves significantly improved density estimation. DDBI captures fine-grained details, while DSBI provides robust performance across all datasets, particularly excelling in modeling complex distributions and mitigating estimation artifacts. Further comparisons across training epochs are available in Table 3.
Energy-based Modeling on MNIST. We applied the proposed $\text{D}^{3}\text{RE}$ framework for density estimation on the MNIST dataset, leveraging pre-trained energy-based models (EBMs) (Choi et al., 2022). Experimental details are in Section C.4. The results are reported in bits-per-dimension (BPD). Results in Table 2 show that $\text{D}^{3}\text{RE}$ achieves the lowest BPD values across all noise types (Gaussian, Copula, and RQ-NSF), outperforming baselines and existing methods. See also Table 3 in the Appendix.
@@ 4.2 Mutual Information Estimation
Mutual information (MI) measures the dependency between two random variables $\mathbf{X}\sim p(\mathbf{x})$ and $\mathbf{Y}\sim q(\mathbf{y})$, defined as $\text{MI}(\mathbf{X},\mathbf{Y})=\mathbb{E}_{p(\mathbf{x},\mathbf{y})}\left[\log\frac{p(\mathbf{x},\mathbf{y})}{p(\mathbf{x})q(\mathbf{y})}\right]$. Here $p(\mathbf{x},\mathbf{y})$ be their joint density. $q(\mathbf{y})=\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{E}_{d})$ and $p(\mathbf{x})=\mathcal{N}(\mathbf{0},\mathbf{E}_{d})$, with $\sigma^{2}=1e-6$ and $d=\{40,80,120\}$, are two $d$-dimensional correlated Gaussian distributions. The experimental setup in DRE-$\infty$ (Choi et al., 2022) is adapted to implement $\text{D}^{3}\text{RE}$. DRE-$\infty$ serves as the benchmark. More details can be found in Section C.3.
The evolution of estimated MI across epochs for $d=\{40,80,120\}$, comparing $\text{D}^{3}\text{RE}$ with DRE-$\infty$, are analyzed. Results in Figure 3 show that the red (DSBI) and green (DDBI) curves outperform the blue and yellow (DRE-$\infty$) curves in two aspects. First, $\text{D}^{3}\text{RE}$ converges to the true MI value more rapidly as it expands trajectory sets (see Corollary 3.3), improving interpolation accuracy. Second, it exhibits greater stability with fewer fluctuations around the true MI, indicating more reliable estimates.
@@ 4.3 Analysis and Discussion
[…]
@@
The ablation study on varying $\gamma^{2}$ values (Figure 5) shows distinct convergence behaviors in MI estimation. For all dimensions ($d=\{40,80,120\}$), smaller $\gamma^{2}$ values ($\leq 0.01$) lead to faster convergence, especially in lower dimensions ($d=40$), but excessively small values ($\gamma^{2}=0.001$) cause instability later. Larger $\gamma^{2}$ values ($\geq 0.1$) converge more slowly but stabilize over time, particularly in higher dimensions ($d=120$). $\gamma^{2}=0.1$ offers a balance between speed and stability across all dimensions, suggesting that moderate regularization provides the best MI estimation performance. More results are provided in Section C.5.
Ablation Study on GD. To evaluate the effectiveness of the proposed GD module, we conduct an ablation study by comparing model performance with and without GD, as shown in Figure 6. Visually, both DDBI and DSBI show clear improvements in density estimation when GD is applied. Without GD, the estimated densities appear blurrier and miss fine structural details, whereas incorporating GD yields sharper and more realistic patterns.
@@
Ablation Study on OTR. We conduct an ablation study to evaluate the role of OTR, comparing models without OTR (baseline, DI), models with OTR (DI+OTR), and models from the $\text{D}^{3}\text{RE}$ framework (DDBI and DSBI). In Figures 2 and 7, DI generates distorted and misaligned intermediate distributions. This shows its limited ability to align with the target distribution. DI+OTR improves alignment but remains suboptimal. Models from the $\text{D}^{3}\text{RE}$ further enhance distribution quality, with DSBI achieving the most precise alignment. This underscores OTR’s crucial role in improving intermediate distributions.
Figure 7: Density estimation results on $\mathsf{checkerboard}$ for different methods during training (see more results in Figures 9, 10, 11, 12, 13, 14, 15 and 16).
Figure 3 compares MI estimation for DDBI and DSBI across dimensions ($d=80,120$). Both outperform baseline methods, but DSBI converges faster and remains closer to the ground truth. The advantage of OTR becomes more pronounced in high dimensions ($d=120$), where DSBI significantly outperforms DDBI in both speed and accuracy.
Overall, OTR improves intermediate distribution alignment. When combined with diffusion bridges and GD, as in DSBI, it enables more accurate density estimation. Further comparisons are presented in Section C.6.
Number of Function Evaluations. We analyze the impact of OTR on NFE, noting that DI and DDBI do not utilize OTR. It shows that applying OTR significantly reduces NFE. Figure 8 compares NFE across four methods in DRE.
@@ 6 Conclusions
In this work, we propose $\text{D}^{3}\text{RE}$, a unified, robust and efficient framework for density ratio estimation. It provides the first framework for directly addressing both the density-chasm and support-chasm problems, enabling uniformly approximated density-ratio estimation (Proposition 3.5). By incorporating diffusion bridges and GD, we construct DDBI, which expands support coverage (Theorem 3.2, Corollary 3.3) and stabilizes the time score near boundaries (Corollary 3.8). Building upon DDBI, OTR is incorporated to derive the DSBI, which offers more efficient and accurate density ratio estimation (Theorem 3.7) by solving the Schrödinger bridge problem (Proposition 3.6). Together, DDBI and DSBI form the core of the $\text{D}^{3}\text{RE}$ framework, enabling uniformly approximated and stable density-ratio estimation (Proposition 3.5). Extensive experiments validate these findings, demonstrating the superior performance of $\text{D}^{3}\text{RE}$ in tasks such as density-ratio estimation on synthetic data, mutual information estimation, and density estimation.
While $\text{D}^{3}\text{RE}$ advances density-ratio estimation methods, several directions remain open. Future work could explore adaptive or learned solvers to reduce function evaluation overhead, as well as more expressive interpolants to further improve robustness in handling complex or multi-modal distributions.
@@ Appendix A Proofs
=$=\{\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{1}\mid\mathbf{x}_{0}\in\mathsf{supp}(q_{0}),\mathbf{x}_{1}\in\mathsf{supp}(q_{1}),\alpha_{t}+\beta_{t}=1\ \text{or}\ \alpha_{t}^{2}+\beta_{t}^{2}=1\},$
@@
[…]
@@
[…]
@@
where $\mathbf{R}_{t}=-\beta_{t}\alpha_{t}^{-1}\nabla_{\mathbf{x}_{t}}\mathbf{x}_{1}$ denote the residual term in the Jacobian decomposition. By A.2 and the definition of the score function of the conditional distribution, $\nabla_{\mathbf{x}_{t}}\log q_{0}(\mathbf{x}_{0}\mid\mathbf{x}_{t})=-\nabla_{\mathbf{x}_{t}}\mathbf{x}_{0}\cdot\nabla_{\mathbf{x}_{0}}\log q_{0}(\mathbf{x}_{0}\mid\mathbf{x}_{t})$, the Lipschitz continuity of the conditional scores implies $\|\nabla_{\mathbf{x}_{t}}\mathbf{x}_{0}\|_{\text{op}}\leq L\alpha_{t}^{-1}$ and $\|\nabla_{\mathbf{x}_{t}}\mathbf{x}_{1}\|_{\text{op}}\leq L\beta_{t}^{-1}$. Here $\|\cdot\|_{\text{op}}$ denotes the operator norm. Hence, the absolute value of the trace of the residual term satisfies:
@@
(3) Lower bound of $\mathbb{E}_{q_{t}}\left[\left|\partial_{t}\log q_{t}\right|\right]$. Finally, bring Equation 33 and Equation 34 back to Equation 28, the lower bound of $\mathbb{E}_{q_{t}}\left[\left|\partial_{t}\log q_{t}\right|\right]$ can be derived:
@@
The residual term $\mathcal{O}(\sqrt{\mathbb{E}_{q_{t}}[\|\nabla\log q_{t}\|^{2}]})$ is finite under A.2 (Lipschitz scores imply bounded Fisher information). Hence, the divergence term dominates asymptotically.
@@
[…]
$=r(\mathbf{x})+\frac{\varepsilon}{2}\underbrace{\left[\frac{\nabla_{\mathbf{x}}^{2}q_{1}(\mathbf{x})}{q_{0}(\mathbf{x})}-r(\mathbf{x})\frac{\nabla_{\mathbf{x}}^{2}q_{0}(\mathbf{x})}{q_{0}(\mathbf{x})}\right]}_{\mathcal{E}(\mathbf{x})}+\mathcal{O}(\varepsilon^{2}).$
To bound $r^{\prime}(\mathbf{x})-r(\mathbf{x})$ in $L^{\infty}$, the supremum can be computed under A.1:
$\|r^{\prime}-r\|_{L^{\infty}}$ $\leq\frac{\varepsilon}{2}\sup_{\mathbf{x}}|\mathcal{E}(\mathbf{x})|+\mathcal{O}(\varepsilon^{2})$ (40)
@@
Under A.1, $\inf\limits_{\mathbf{x}}q_{0}(\mathbf{x})$ has lower bound and these norms have upper bound. Then as $\varepsilon\to 0$,
@@
Based on Theorem 2.4 of (Léonard, 2014), (De Bortoli et al., 2021) established that the solution to the SB problem as detailed in Equation 19, $\pi^{\star}$, satisfies the SB conditions: (1) the optimization problem for $\pi^{\star}$ is equivalent to the entropically regularied OT problem, with the optimal coupling $\pi_{2\gamma^{2}}$ defined in Equation 17; (2) for samples $\left(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime}\right)\sim\pi^{\star}$, the associated conditional path distributions $\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ minimize the KL divergence: $\mathbb{E}_{(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}}\mathrm{KL}\left(\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\|\pi_{\mathrm{ref}}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\right)$, where $\pi_{\mathrm{ref}}$ is the reference path distribution satisfying $\log\pi_{\mathrm{ref}}(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\frac{\|\mathbf{x}_{0}^{\prime}-\mathbf{x}_{1}^{\prime}\|^{2}}{2\gamma^{2}}+\text{const}$. These conditional distributions are optimized using Brownian bridges of diffusion scale $\gamma$, conditioned on the endpoints $\mathbf{x}_{0}^{\prime}$ and $\mathbf{x}_{1}^{\prime}$. The marginal distribution at intermediate time $t$ along the Brownian bridge is given by $q_{t}(\mathbf{x}\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\mathcal{N}(\mathbf{x}\mid(1-t)\mathbf{x}_{0}^{\prime}+t\mathbf{x}_{1}^{\prime},t(1-t)\gamma^{2}\mathbf{E}_{d})$ (Tong et al., 2024). Since our proposed OTR method uses Sinkhorn’s algorithm to solve the entropically regularized OT problem, and the probability paths of our DDBI with $\alpha_{t}=1-t$ and $\beta_{t}=t$ align with those of the Brownian bridge, a trajectory $\mathbf{x}_{t}^{\prime}$ generated by first sampling $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}$, then sampling $\mathbf{x}_{t}\sim q_{t}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ satisfies the SB conditions, thus verifying the proposition. ∎

### v5

@@ 4.1 Density Estimation
2-D Synthetic Datasets. We trained DRE-$\infty$ (baseline) and $\text{D}^{3}\text{RE}$ (ours) on eight 2-D synthetic datasets, including $\mathsf{swissroll}$, $\mathsf{circles}$, $\mathsf{rings}$, $\mathsf{moons}$, $\mathsf{8gaussians}$, $\mathsf{pinwheel}$, $\mathsf{2spirals}$, and $\mathsf{checkerboard}$, for 20,000 epochs using the joint score matching loss (details in Sec. C.2). The density estimation results are shown in Fig. 2.
@@
In contrast, the $\text{D}^{3}\text{RE}$ framework, leveraging both DDBI and DSBI, achieves significantly improved density estimation. DDBI captures fine-grained details, while DSBI provides robust performance across all datasets, particularly excelling in modeling complex distributions and mitigating estimation artifacts. Further comparisons across training epochs are available in Tab. 3.
Energy-based Modeling on MNIST. We applied the proposed $\text{D}^{3}\text{RE}$ framework for density estimation on the MNIST dataset, leveraging pre-trained energy-based models (EBMs) (Choi et al., 2022). Experimental details are in Sec. C.4. The results are reported in bits-per-dimension (BPD). Results in Tab. 2 show that $\text{D}^{3}\text{RE}$ achieves the lowest BPD values across all noise types (Gaussian, Copula, and RQ-NSF), outperforming baselines and existing methods. See also Tab. 3 in the Appendix.
@@ 4.2 Mutual Information Estimation
Mutual information (MI) measures the dependency between two random variables $\mathbf{X}\sim p(\mathbf{x})$ and $\mathbf{Y}\sim q(\mathbf{y})$, defined as $\text{MI}(\mathbf{X},\mathbf{Y})=\mathbb{E}_{p(\mathbf{x},\mathbf{y})}\left[\log\frac{p(\mathbf{x},\mathbf{y})}{p(\mathbf{x})q(\mathbf{y})}\right]$. Here $p(\mathbf{x},\mathbf{y})$ be their joint density. $q(\mathbf{y})=\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{E}_{d})$ and $p(\mathbf{x})=\mathcal{N}(\mathbf{0},\mathbf{E}_{d})$, with $\sigma^{2}=1e-6$ and $d=\{40,80,120\}$, are two $d$-dimensional correlated Gaussian distributions. The experimental setup in DRE-$\infty$ (Choi et al., 2022) is adapted to implement $\text{D}^{3}\text{RE}$. DRE-$\infty$ serves as the benchmark. More details can be found in Sec. C.3.
The evolution of estimated MI across epochs for $d=\{40,80,120\}$, comparing $\text{D}^{3}\text{RE}$ with DRE-$\infty$, are analyzed. Results in Fig. 3 show that the red (DSBI) and green (DDBI) curves outperform the blue and yellow (DRE-$\infty$) curves in two aspects. First, $\text{D}^{3}\text{RE}$ converges to the true MI value more rapidly as it expands trajectory sets (see Corollary 3.3), improving interpolation accuracy. Second, it exhibits greater stability with fewer fluctuations around the true MI, indicating more reliable estimates.
@@ 4.3 Analysis and Discussion
Ablation Study on $\gamma^{2}$. The ablation study on $\gamma^{2}$ for density estimation (Fig. 4) reveals systematic trade-offs in performance across regularization strengths. For small $\gamma^{2}=0.001$, the model achieves rapid initial alignment with the ground truth distribution (first row) but exhibits overfitting artifacts in later epochs, manifesting as irregular density peaks and deviations from the smooth ground truth structure. Intermediate values ($\gamma^{2}=0.01$$0.1$) demonstrate balanced behavior: $\gamma^{2}=0.01$ preserves finer details while maintaining stability, and $\gamma^{2}=0.1$ produces smoother approximations with minimal divergence from the true distribution. Larger $\gamma^{2}$ values ($\geq 0.5$) induce excessive regularization, leading to oversmoothed estimates that fail to capture critical modes of the 2-D data, particularly in high-density regions. Notably, $\gamma^{2}=0.1$ achieves the closest visual and structural resemblance to the ground truth, suggesting its suitability for low-dimensional tasks requiring both fidelity and robustness. These results underscore the necessity of tuning $\gamma^{2}$ to mitigate under-regularization artifacts while preserving distributional complexity.
@@
The ablation study on varying $\gamma^{2}$ values (Fig. 5) shows distinct convergence behaviors in MI estimation. For all dimensions ($d=\{40,80,120\}$), smaller $\gamma^{2}$ values ($\leq 0.01$) lead to faster convergence, especially in lower dimensions ($d=40$), but excessively small values ($\gamma^{2}=0.001$) cause instability later. Larger $\gamma^{2}$ values ($\geq 0.1$) converge more slowly but stabilize over time, particularly in higher dimensions ($d=120$). $\gamma^{2}=0.1$ offers a balance between speed and stability across all dimensions, suggesting that moderate regularization provides the best MI estimation performance. More results are provided in Sec. C.5.
Ablation Study on GD. To evaluate the effectiveness of the proposed GD module, we conduct an ablation study by comparing model performance with and without GD, as shown in Fig. 6. Visually, both DDBI and DSBI show clear improvements in density estimation when GD is applied. Without GD, the estimated densities appear blurrier and miss fine structural details, whereas incorporating GD yields sharper and more realistic patterns.
@@
Ablation Study on OTR. We conduct an ablation study to evaluate the role of OTR, comparing models without OTR (baseline, DI), models with OTR (DI+OTR), and models from the $\text{D}^{3}\text{RE}$ framework (DDBI and DSBI). In Figs. 2 and 7, DI generates distorted and misaligned intermediate distributions. This shows its limited ability to align with the target distribution. DI+OTR improves alignment but remains suboptimal. Models from the $\text{D}^{3}\text{RE}$ further enhance distribution quality, with DSBI achieving the most precise alignment. This underscores OTR’s crucial role in improving intermediate distributions.
Figure 7: Density estimation results on $\mathsf{checkerboard}$ for different methods during training (see more results in Figs. 9, 10, 11, 12, 13, 14, 15 and 16).
Fig. 3 compares MI estimation for DDBI and DSBI across dimensions ($d=80,120$). Both outperform baseline methods, but DSBI converges faster and remains closer to the ground truth. The advantage of OTR becomes more pronounced in high dimensions ($d=120$), where DSBI significantly outperforms DDBI in both speed and accuracy.
Overall, OTR improves intermediate distribution alignment. When combined with diffusion bridges and GD, as in DSBI, it enables more accurate density estimation. Further comparisons are presented in Sec. C.6.
Number of Function Evaluations. We analyze the impact of OTR on NFE, noting that DI and DDBI do not utilize OTR. It shows that applying OTR significantly reduces NFE. Fig. 8 compares NFE across four methods in DRE.
@@ 6 Conclusions
In this work, we propose $\text{D}^{3}\text{RE}$, a unified, robust and efficient framework for density ratio estimation. It provides the first framework for directly addressing both the density-chasm and support-chasm problems, enabling uniformly approximated density-ratio estimation (Proposition 3.5). By incorporating diffusion bridges and GD, we construct DDBI, which expands support coverage (Theorem 3.2, Corollary 3.3) and stabilizes the time score near boundaries (Corollary 3.8). Building upon DDBI, OTR is incorporated to derive the DSBI, which offers more efficient and stable density ratio estimation (Theorem 3.7) by solving the Schrödinger bridge problem (Proposition 3.6). Together, DDBI and DSBI form the core of the $\text{D}^{3}\text{RE}$ framework, enabling uniformly approximated and stable density-ratio estimation (Proposition 3.5). Extensive experiments validate these findings, demonstrating the superior performance of $\text{D}^{3}\text{RE}$ in tasks such as density-ratio estimation on synthetic data, mutual information estimation, and density estimation.
While $\text{D}^{3}\text{RE}$ advances density ratio estimation methods, several directions remain open. Future work could explore adaptive or learned solvers to reduce function evaluation overhead, as well as more expressive interpolants to further improve robustness in handling complex or multi-modal distributions.
@@ Appendix A Proofs
$=\{\alpha_{t}\mathbf{x}_{0}+\beta_{t}\mathbf{x}_{1}\mid\mathbf{x}_{0}\in\mathsf{supp}(q_{0}),\mathbf{x}_{1}\in\mathsf{supp}(q_{1}),\alpha_{t}+\beta_{t}=1\ \text{or}\ \alpha_{t}^{2}+\beta_{t}^{2}=1\},$
@@
where $\nabla\cdot$ and $\nabla$ are the divergence and gradient operators w.r.t. $\mathbf{x}_{t}$. $\mathbf{u}_{t}(\mathbf{x}_{t})=\mathbb{E}_{\pi}[\dot{\alpha}_{t}\mathbf{X}_{0}+\dot{\beta}_{t}\mathbf{X}_{1}\mid\mathbf{X}_{t}=\mathbf{x}_{t}]$ is the drift term. Taking expectations over $q_{t}$ and applying the triangle inequality and Cauchy-Schwarz inequality to this equation:
@@
$\geq\mathbb{E}_{q_{t}}[\left|\nabla\cdot\mathbf{u}_{t}\right|]-\mathbb{E}_{q_{t}}[\|\mathbf{u}_{t}\|\cdot\|\nabla\log q_{t}\|]\quad\left(\text{Cauchy-Schwarz inequality}\right)$
@@
where $\mathbf{R}_{t}=-\beta_{t}\alpha_{t}^{-1}\nabla_{\mathbf{x}_{t}}\mathbf{x}_{1}$ denote the residual term in the Jacobian decomposition. By Assumption A.2 and the definition of the score function of the conditional distribution, $\nabla_{\mathbf{x}_{t}}\log q_{0}(\mathbf{x}_{0}\mid\mathbf{x}_{t})=-\nabla_{\mathbf{x}_{t}}\mathbf{x}_{0}\cdot\nabla_{\mathbf{x}_{0}}\log q_{0}(\mathbf{x}_{0}\mid\mathbf{x}_{t})$, the Lipschitz continuity of the conditional scores implies $\|\nabla_{\mathbf{x}_{t}}\mathbf{x}_{0}\|_{\text{op}}\leq L\alpha_{t}^{-1}$ and $\|\nabla_{\mathbf{x}_{t}}\mathbf{x}_{1}\|_{\text{op}}\leq L\beta_{t}^{-1}$. Here $\|\cdot\|_{\text{op}}$ denotes the operator norm. Hence, the absolute value of the trace of the residual term satisfies:
@@
(3) Lower bound of $\mathbb{E}_{q_{t}}\left[\left|\partial_{t}\log q_{t}\right|\right]$. Finally, bring Eq. 34 and Eq. 35 back to Eq. 29, the lower bound of $\mathbb{E}_{q_{t}}\left[\left|\partial_{t}\log q_{t}\right|\right]$ can be derived:
@@
The residual term $\mathcal{O}(\sqrt{\mathbb{E}_{q_{t}}[\|\nabla\log q_{t}\|^{2}]})$ is finite under Assumption A.2 (Lipschitz scores imply bounded Fisher information). Hence, the divergence term dominates asymptotically.
@@
$=\frac{q_{1}(\mathbf{x})}{q_{0}(\mathbf{x})}+\frac{\varepsilon}{2}\frac{\nabla_{\mathbf{x}}^{2}q_{1}(\mathbf{x})}{q_{0}(\mathbf{x})}-\frac{\varepsilon}{2}r(\mathbf{x})\frac{\nabla_{\mathbf{x}}^{2}q_{0}(\mathbf{x})}{q_{0}(\mathbf{x})}+\mathcal{O}(\varepsilon^{2})\quad(\text{First-order expansion of a fraction})$
$=r(\mathbf{x})+\frac{\varepsilon}{2}\underbrace{\left[\frac{\nabla_{\mathbf{x}}^{2}q_{1}(\mathbf{x})}{q_{0}(\mathbf{x})}-r(\mathbf{x})\frac{\nabla_{\mathbf{x}}^{2}q_{0}(\mathbf{x})}{q_{0}(\mathbf{x})}\right]}_{\Delta(\mathbf{x})}+\mathcal{O}(\varepsilon^{2}).$
To bound $r^{\prime}(\mathbf{x})-r(\mathbf{x})$ in $L^{\infty}$, the supremum can be computed under Assumption A.1:
$\|r^{\prime}-r\|_{L^{\infty}}$ $\leq\frac{\varepsilon}{2}\sup_{\mathbf{x}}|\Delta(\mathbf{x})|+\mathcal{O}(\varepsilon^{2})$ (41)
@@
Under Assumption A.1, $\inf\limits_{\mathbf{x}}q_{0}(\mathbf{x})$ has lower bound and these norms have upper bound. Then as $\varepsilon\to 0$,
@@
Based on Theorem 2.4 of (Léonard, 2014), (De Bortoli et al., 2021) established that the solution to the SB problem as detailed in Eq. 19, $\pi^{\star}$, satisfies the SB conditions: (1) the optimization problem for $\pi^{\star}$ is equivalent to the entropically regularied OT problem, with the optimal coupling $\pi_{2\gamma^{2}}$ defined in Eq. 17; (2) for samples $\left(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime}\right)\sim\pi^{\star}$, the associated conditional path distributions $\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ minimize the KL divergence: $\mathbb{E}_{(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}}\mathrm{KL}\left(\pi^{\star}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\|\pi_{\mathrm{ref}}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\right)$, where $\pi_{\mathrm{ref}}$ is the reference path distribution satisfying $\log\pi_{\mathrm{ref}}(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\frac{\|\mathbf{x}_{0}^{\prime}-\mathbf{x}_{1}^{\prime}\|^{2}}{2\gamma^{2}}+\text{const}$. These conditional distributions are optimized using Brownian bridges of diffusion scale $\gamma$, conditioned on the endpoints $\mathbf{x}_{0}^{\prime}$ and $\mathbf{x}_{1}^{\prime}$. The marginal distribution at intermediate time $t$ along the Brownian bridge is given by $q_{t}(\mathbf{x}\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})=\mathcal{N}(\mathbf{x}\mid(1-t)\mathbf{x}_{0}^{\prime}+t\mathbf{x}_{1}^{\prime},t(1-t)\gamma^{2}\mathbf{E}_{d})$ (Tong et al., 2024). Since our proposed OTR method uses Sinkhorn’s algorithm to solve the entropically regularized OT problem, and the probability paths of our DDBI with $\alpha_{t}=1-t$ and $\beta_{t}=t$ align with those of the Brownian bridge, a trajectory $\mathbf{x}_{t}^{\prime}$ generated by first sampling $(\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})\sim\pi^{\star}$, then sampling $\mathbf{x}_{t}\sim q_{t}(\cdot\mid\mathbf{x}_{0}^{\prime},\mathbf{x}_{1}^{\prime})$ satisfies the SB conditions, thus verifying the proposition. ∎